# Ore extensions of multiplier Hopf coquasigroups

### v1

@@
Abstract In this paper, Ore extensions of multiplier Hopf coquasigroups are studied. Necessary and sufficient conditions for the Ore extension of a multiplier Hopf coquasigroup to be a multiplier Hopf coquasigroup are given. Then the isomorphism between two Ore extensions is discussed.
@@ 1 Introduction
The Ore extension is a well-known method of ring extensions to construct noncommutative rings and algebras. In Hopf algebra theory, Ore extension were also served as a key tool to create noncommutative and non-cocommutative Hopf algebras, such as Panov [9], Beattie [1, 2] and Nenciu [7, 8].
Over time, Hopf algebras has been evolved into broader frameworks. In the 1990s, Van Daele [10] pioneered the theory of multiplier Hopf algebras, breaking the unit constraint of traditional Hopf algebras and offering new tools for quantum groups and noncommutative algebra. Subsequently, its theory was gradually expanded to duality theory, representation theory, and quantum applications. Meanwhile, the theory of Hopf coquasigroups advanced through the groundbreaking work of Klim and Majid [6]. They introduced weak coassociativity to resolve dual construction challenges in non-coassociative structures. Their work also revealed deep connections with algebraic topology.
Subsequently, some algebraists investigated the feasibility of Ore extensions on these structures (Figure 1). Zhao and Lu [14] initiated research on Ore extensions of multiplier Hopf algebras, while Jiao and Meng [3] extended this to Hopf coquasigroups, proposing Hopf coquasigroup-Ore extensions.
[…]
@@
Based on these advances, a natural question arises: Could Ore extensions also hold for multiplier Hopf coquasigroups? This is the main motivation.
This paper is organized as follows. In section 2, we introduce some basic notions: Ore extension, multiplier algebras and multiplier Hopf coquasigroups, which will be used in the following sections.
In section 3, we present the concept of Ore extensions in the context of multiplier Hopf coquasigroups. Moreover, we present the main theorem of this paper (Theorem 3.4), which gives a necessary and sufficient condition for Ore extensions of regular multiplier Hopf coquasigroups to be regular multiplier Hopf coquasigroups. This theorem gives a positive answer to the above question.
In section 4, we give the notion of isomorphism for Ore extensions, and show sufficient conditions for two such extensions to be isomorphic.
@@ 2 Preliminaries
Throughout this paper, all linear spaces we considered are over a fixed field $k$ (e.g., the complex number field $\mathds{C}$).
@@ 2.1 Multiplier Hopf coquasigroups
Let $A$ be an (associative) algebra over $k$, with or without identity, but the product is nondegenerate. This means that, whenever $a\in A$ and $ab=0$ for all $b\in A$ or $ba=0$ for all $b\in A$, we must have that $a=0$.
Recall from [10, 11] that a left multiplier of $A$ is a linear map $\rho:A\rightarrow A$ such that $\rho(ab)=\rho(a)b$ for all $a,b\in A$. Similarly, a right multiplier of $A$ is a linear map $\rho:A\rightarrow A$ such that $\rho(ab)=a\rho(b)$ for all $a,b\in A$. A multiplier of $A$ is a pair $(\rho_{1},\rho_{2})$ of a left and a right multiplier such that $\rho_{2}(a)b=a\rho_{1}(b)$ for all $a,b\in A$. For any algebra $A$, there are a left multiplier algebra, a right multiplier algebra, and a multiplier algebra of $A$, which are respectively denoted by $L(A)$, $R(A)$ and $M(A)$. In particular, $M(A)$ is characterized as the largest algebra with identity containing $A$ as an essential two-sided ideal. Additionally, we still have that, whenever $a\in M(A)$ and $ab=0$ for all $b\in A$ or $ba=0$ for all $b\in A$, again $a=0$. Furthermore, we consider the tensor algebra $A\otimes A$. It is still nondegenerate and we have its multiplier algebra $M(A\otimes A)$. There are natural imbeddings
@@
In general, when $A$ has no identity, these two inclusions are strict. If $A$ already has an identity $1_{A}$, the product is obviously nondegenerate and $M(A)=A$ and $M(A\otimes A)=A\otimes A$.
Let $A$ and $B$ be nondegenerate algebras, if homomorphism $f:A\longrightarrow M(B)$ is nondegenerate (i.e., $f(A)B=B$ and $Bf(A)=B$), then it has a unique extension to a homomorphism $M(A)\longrightarrow M(B)$, we also denote it $f$.
A regular multiplier Hopf coquasigroup introduced in [13] is a nondegenerate associative algebra $A$ equipped with algebra homomorphisms $\Delta:A\longrightarrow M(A\otimes A)$(coproduct), $\varepsilon:A\longrightarrow k$ (counit), and a linear map $S:A\longrightarrow A$ (antipode) such that
@@
    $T_{1}(a\otimes b)=\Delta(a)(1\otimes b)$ and $T_{2}(a\otimes b)=(a\otimes 1)\Delta(b)$ belong to $A\otimes A$ for any $a,b\in A$.
@@
    The counit satisfies $(\varepsilon\otimes id)T_{1}(a\otimes b)=ab=(id\otimes\varepsilon)T_{2}(a\otimes b)$.
@@
    $S$ is antimultiplicative and anticomultiplicative such that for any $a,b\in A$
@@
Multiplier Hopf coquasigroup $(A,\Delta)$ is called regular, if the antipode $S$ is bijective. $T_{1}$ and $T_{2}$ in the above are usually called Galois maps.
Multiplier Hopf coquasigroups introduced in[13] generalized Hopf coquasigroups to a non-unital case. Unlike Hopf coquasigroups in [5, 6], the underlying algebra of multiplier Hopf coquasigroup $A$ need not have an identity in general, and the comultiplication $\Delta$ does not map $A$ into $A\otimes A$ but rather into its multiplier algebra $M(A\otimes A)$. Multiplier Hopf coquasigroups weaken the coassociativity of multiplier Hopf algebra, and unify the notions of multiplier Hopf algebras and Hopf coquasigroups.
We use the adapted Sweedler notation introduced in [12] for the comultiplication: $\Delta(a)=a_{(1)}\otimes a_{(2)}$ for $a\in A$.
2.2 Ore extention
@@
First, let us recall the definition of Ore extension of an algebra from [4]. Let $A$ be an algebra, and $\tau{}$ be an algebra endomorphism of $A$. A linear endomorphism $\delta{}$ of $A$ is called an $\tau$-derivation $\delta$ of $A$ if for all $a,b\in A$,
$\delta(ab)=\delta(a)b+\tau(a)\delta(b)$
The above condition implies $\delta(1_{A})=0$.
The Ore extension $A[y;\tau,\delta]$ of an algebra $A$ is an algebra generated by the algebra $A$ with the relation
@@
Ore extension of Hopf algebra
Let $A$ and $A[y;\tau,\delta]$ be Hopf algebras over $k$. The Hopf algebra $A[y;\tau,\delta]$ is called the Hopf-Ore extension if $\Delta(y)=y\otimes 1+r\otimes y$ for some $r\in A$ and $A$ is a Hopf subalgebra in $A[y;\tau,\delta]$.
Following Theorem 1.3 in [9], the necessary and sufficient conditions for the Hopf algebra $A[y;\tau,\delta]$ to be a Hopf-Ore extension are given as follows.
(1) there is a character $\chi:A\rightarrow k$ such that
@@
(2) the following relation holds:
@@
(3) the $\tau$-derivation $\delta$ satisfies the relation
@@
Let $A$ and $A[y;\tau,\delta]$ be Hopf coquasigroups, recall from [3] that the Hopf coquasigroup $A[y;\tau,\delta]$ is a Hopf coquasigroups-Ore extension if $\Delta(y)=y\otimes 1+r\otimes y$ for some $r\in A$ and $A$ is a sub-Hopf coquasigroup of $A[y;\tau,\delta]$.
From Theorem 3.3 in [3], for a Hopf coquasigroup $A$, the Hopf coquasigroup $H=A[y,\tau,\delta]$ becomes an Hopf coquasigroup-Ore extension if and only if the following conditions hold:
(1) There exists a character $\chi:A\rightarrow k$, such that for every $a\in A$, satisfying
$\alpha(a)=\chi(a_{(1)})a_{(2)}.$
(2) The subsequent relations hold:
@@
(3) For the $\alpha$-derivation $\delta$, the relation
@@
Ore extension of Multiplier Hopf algebras
Let $(A,\Delta)$ be a regular multiplier Hopf algebra and $A[y;\tau,\delta]$ be an Ore extension of the algebra $A$.
If $\tau$ is a nondegenerate algebra endomorphism, then $\tau$-derivation $\delta$ uniquely extends to the multiplier algebra $M(A)$ (i.e., $\overline{\delta}:M(A)\rightarrow M(A)$). If $\tau$ is surjective, following Proposition 2.3 of Section 2 in [14], the product of $A[y,\tau,\delta]$ is nondegenerate. There are some natural extensions: the comultiplication $\overline{\Delta}:M(A)[y;\tau,\delta]\rightarrow M(A[y;\tau,\delta]\otimes A[y;\tau,\delta])$, the antipode $\overline{S}:M(A)[y;\tau,\delta]\rightarrow M(A[y;\tau,\delta])$, the counit $\overline{\varepsilon}:M(A)[y;\tau,\delta]\rightarrow k$.
If restricting these extensions to the subspace $A[y;\tau,\delta]$, i.e., $\overline{\Delta}\big{|}_{A[y;\tau,\delta]}:A[y;\tau,\delta]\rightarrow M(A[y;\tau,\delta]\otimes A[y;\tau,\delta])$, $\overline{S}\big{|}_{A[y;\tau,\delta]}:A[y;\tau,\delta]\rightarrow M(A[y;\tau,\delta])$, $\overline{\varepsilon}\big{|}_{A[y;\tau,\delta]}:A[y;\tau,\delta]\rightarrow k$, we get the comultiplication $\Delta^{\prime}$, counit $\varepsilon^{\prime}$, and the antipode $S^{\prime}$ on the Ore extension $(A[y;\tau,\delta],\Delta^{\prime})$ of $A$.
$(A[y;\tau,\delta],\Delta^{\prime})$ is called a Hopf-Ore extension of the multiplier Hopf algebra $A$, if $(A[y;\tau,\delta],\Delta^{\prime})$ is a regular multiplier Hopf algebra such that $\overline{\Delta}(y)=y\otimes 1+r\otimes y$ for some $r\in M(A)$.
For a regular multiplier Hopf algebra $A$ with Ore extension $A[y;\tau,\delta]$, where $\tau$ is surjective. $A[y;\tau,\delta]$ constitutes a Hopf-Ore extension if and only if there exists a group-like element $r\in M(A)$ satisfying the following equivalent conditions:
$\tau(a)=(\chi\otimes id)\Delta(a)=r\big{(}(id\otimes\chi)\Delta(a)\big{)}r^{-1}.$
(2) For the $\tau$-derivation $\delta$, for all $a\in A$, the relation
$\Delta\big{(}\delta(a)\big{)}=(\delta\otimes id)\Delta(a)+(r\otimes 1)\big{(}(id\otimes\delta)\Delta(a)\big{)}$
is satisfied.
As introduced in Section 2.1 of this paper, in the following, we also denote these unique extensions $\overline{\Delta}$, $\overline{S}$, $\overline{\varepsilon}$ as $\Delta$, $S$, $\varepsilon$ without confusion.
@@ 3 Ore extensions of multiplier Hopf coquasigroups
Let $(A,\Delta)$ be a regular multiplier Hopf coquasigroup introduced in [13]. The aim of this section is to show a criterion for an Ore extension of a regular multiplier Hopf coquasigroup to be a regular multiplier Hopf coquasigroup.
@@
First, let us introduce the definition of Ore extension of regular multiplier Hopf coquasigroups, which genelized the Ore extension of Hopf coquasigroups and multiplier Hopf algebras introduced in [3, 14].
@@
Definition 3.1. Let $A$ and $A[y;\tau,\delta]$ be regular multiplier Hopf coquasigroups. The multiplier Hopf coquasigroup $(A[y;\tau,\delta],\Delta)$ is called the Ore extension of multiplier Hopf coquasigroups (or MHC-Ore extension in short) if $\Delta(y)=y\otimes r_{1}+r_{2}\otimes y$ for some $r_{1},r_{2}\in M(A)$.
Remark 3.2. (1) As in [14], $y\otimes 1+r\otimes y$ in above definition belongs to $M(A[y;\tau,\delta]\otimes A[y;\tau,\delta])$.
(2) If the coproduct $\Delta$ of multiplier Hopf coquasigroup $A$ is coassociative, then $A$ is a multiplier Hopf algebra and the Ore extension is actually the one introduced in [14].
@@
(3) If the underlying algebra of multiplier Hopf coquasigroup $A$ has an identity, then $A$ is a Hopf coquasigroup and the Ore extension is actually the one introduced in [3].
By the definition of regular multiplier Hopf coquasigroup, for $y\in A[y;\tau,\delta]$, we have
$(m\otimes\iota)(S\otimes\Delta)\Delta(y)=1_{M(A)}\otimes y=(m\otimes\iota)(\iota\otimes S\otimes\iota)(\iota\otimes\Delta)\Delta(y),$
$(\iota\otimes m)(\Delta\otimes S)\Delta(y)=y\otimes 1_{M(A)}=(\iota\otimes m)(\iota\otimes S\otimes\iota)(\Delta\otimes\iota)\Delta(y),$
that is
$S(y)r_{1(1)}\otimes r_{1(2)}+S(r_{2})y\otimes r_{1}+S(r_{2})r_{2}\otimes y=1_{M(A)}\otimes y$
$=yS(r_{1(1)})\otimes r_{1(2)}+r_{2}S(y)\otimes r_{1}+r_{2}S(r_{2})\otimes y,$
$y\otimes S(r_{1})r_{1}+r_{2}\otimes S(y)r_{1}+r_{2(1)}\otimes S(r_{2(2)})y=y\otimes 1_{M(A)}$
$=y\otimes r_{1}S(r_{1})+r_{2}\otimes yS(r_{1})+r_{2(1)}\otimes r_{2(2)}S(y).$
Then
$\hskip 40.00006pt{r_{1}}S({r_{1}})=S({r_{1}}){r_{1}}=1,\qquad{r_{2}}S({r_{2}})=S({r_{2}}){r_{2}}=1,$
$S(y){r_{1(1)}}\otimes{r_{1(2)}}+S({r_{2}})y\otimes{r_{1}}=yS({r_{1(1)}})\otimes{r_{1(2)}}+{r_{2}}S(y)\otimes{r_{1}}=0,$
${r_{2}}\otimes S(y){r_{1}}+{r_{2(1)}}\otimes S({r_{2(2)}})y={r_{2}}\otimes yS({r_{1}})+{r_{2(1)}}\otimes{r_{2(2)}}S(y)=0.$
Let $r_{1}$ and $r_{2}$ be group-like elements in $M(A)$, then we have
$\Delta(r_{i})=r_{i}\otimes r_{i},\quad\varepsilon(r_{i})=1,\quad S(r_{i})=r_{i}^{-1},\quad i=1,2,$
$\big{(}S(y)r_{1}+S(r_{2})y\big{)}\otimes r_{1}=\big{(}yS(r_{1})+r_{2}S(y)\big{)}\otimes r_{1}=0,$
$r_{2}\otimes\big{(}S(y)r_{1}+S(r_{2})y\big{)}=r_{2}\otimes\big{(}yS(r_{1})+r_{2}S(y)\big{)}=0.$
By substituting the generator $y$ with $y^{\prime}=yr_{1}^{-1}$, it follows that the coproduct map satisfies $\Delta(y^{\prime})=y^{\prime}\otimes 1+r\otimes y^{\prime}$, where $r=r_{2}r_{1}^{-1}$ is a group-like element in the multiplier algebra $M(A)$. In the subsequent discussion, we consistently assume that the element $y$ in the Ore extension of a regular multiplier Hopf coquasigroup, the following relation holds:
$\Delta\left(y\right)=y\otimes 1+r\otimes y$ (3.1)
for some group-like element $r\in M(A)$. As usual, $A{d_{r}}\left(a\right)=raS\left(r\right)=ra{r^{-1}}$.
@@
Now we give the following properties for the Ore extension of a regular multiplier Hopf coquasigroup.
Proposition 3.3. When $A[y;\tau,\delta]$ is a MHC-Ore extension of $A$, then
$\varepsilon(y)=0,\quad S(y)=-r^{-1}y,\quad S^{-1}(y)=-\tau(r)^{-1}y,$
$\hskip 60.00009pt\delta(r)=0,\quad\delta(r^{-1})=0.$
proof. By applying the main property of counit, where $a=a_{(1)}\varepsilon(a_{(2)})$, it follows that $y=\varepsilon(y)+\varepsilon(r)y=\varepsilon(y)+y$, this implies $\varepsilon(y)=0$.
By applying the main property of antipode, where $\varepsilon(a)=a_{(1)}S(a_{(2)})$, it follows that $0=\varepsilon(y)=y+rS(y)$, this implies $S(y)=-r^{-1}y$. Since
$y=S^{-1}S(y)=S^{-1}(-r^{-1}y)=-S^{-1}(r^{-1}y)=-S^{-1}(y)S^{-1}(r^{-1})=-S^{-1}(y)r,$
there is an element $a\in A$ such that $S^{-1}(y)=ay$. And applying the preceding formula we get
$y=-ayr=-a\tau(r)y-a\delta(r),$
this implies that $a=-\tau(r)^{-1}$ and $\delta(r)=0$, so we get that $S^{-1}(y)=-\tau(r)^{-1}y$. Since $\delta$ is a $\tau-derivation$, we have
$0=\delta(1)=\delta(rr^{-1})=\delta(r)r^{-1}+\tau(r)\delta(r^{-1}),$

### v2

@@
Abstract In this paper, Ore extensions of multiplier Hopf coquasigroups are studied. Necessary and sufficient conditions for the Ore extension of a regular multiplier Hopf coquasigroup to be a multiplier Hopf coquasigroup are given. Furthermore, the isomorphism between two such Ore extensions is discussed.
@@ 1 Introduction
Hopf algebras have evolved into increasingly general formalisms over the past decades. In the early 1990s, Van Daele [10] introduced the concept of multiplier Hopf algebras, a framework that dispenses with the unitality requirement intrinsic to classical Hopf algebras and thereby furnishes new tools for the study of quantum groups and non-commutative algebras. This theory has since been systematically expanded to include an intrinsic duality theory, a comprehensive representation theory, and applications in quantum field theory. Hopf coquasigroups have undergone substantial development, particularly through the foundational work in 2010 of Klim and Majid [6], who formulated a weak form of coassociativity. Their approach establishes profound links with algebraic topology and opens new avenues for the study of non-(co)associative symmetries.
The Ore extension is a well-known method in ring extensions used to construct noncommutative rings and algebras. Moreover, within the theory of Hopf algebras, Ore extensions are significant for constructing noncommutative and non-cocommutative Hopf algebras, as demonstrated in the works of Panov [9], Beattie [1, 2], and Nenciu [7, 8]. Subsequently, researchers investigated the feasibility of endowing these generalized Hopf-type structures with Ore extensions (see Figure 1). Zhao and Lu [16] initiated the systematic study of Ore extensions in the context of multiplier Hopf algebras, while Jiao and Meng [3] formulated Ore extensions for Hopf coquasigroups and established the notion of Hopf coquasigroup–Ore extensions.
Ore Extensions ofHopf algebrasOre Extensions ofHopf coquasigroupsOre Extensions ofmultiplier Hopf algebrasOre Extensions ofmultiplier Hopf coquasigroupsJiao & Meng[9]Zhao & Lu[16] ??
@@
In 2022, Yang [14] established a unifying framework for multiplier Hopf algebras and Hopf coquasigroups through the introduction of multiplier Hopf coquasigroups. This theoretical advancement not only bridged these two important algebraic structures but also provided a solution to the biduality problem for a class of infinite-dimensional Hopf quasigroups. Motivated by these recent advances, the present work addresses the following fundamental question: Can Ore extensions be realized within the framework of multiplier Hopf coquasigroups?
This paper provides a positive answer, and its structure is as follows. In Section 2, we provide an introduction to key concepts: Ore extensions, multiplier algebras, and multiplier Hopf coquasigroups, which will be referenced in subsequent sections.
Section 3 discusses Ore extensions within the framework of multiplier Hopf coquasigroups. Additionally, one of the main results of this study (Theorem 3.4) is presented, offering a definitive condition for Ore extensions of regular multiplier Hopf coquasigroups to also be regular. This theorem affirms the previous question. Meanwhile, we consider the *-algebra version.
Moving on to Section 4, we discuss the isomorphism for Ore extensions and establish the criteria under which two such extensions are considered isomorphic.
@@ 2 Preliminaries
Throughout this paper, all linear spaces under consideration are defined over a fixed field $k$ (for example, the field of complex numbers $\mathds{C}$).
@@ 2.1 Multiplier Hopf coquasigroups
Let $A$ be an (associative) algebra over $k$, with or without an identity element, but with a nondegenerate product. This condition implies that if $a\in A$ and $ab=0$ for all $b\in A$ or $ba=0$ for all $b\in A$, then $a$ must be equal to 0.
Recall from [10, 11, 13] that a left multiplier of $A$ is a linear map $\rho_{1}:A\rightarrow A$ satisfying $\rho_{1}(ab)=\rho_{1}(a)b$ for all $a,b\in A$. Similarly, a right multiplier of $A$ is a linear map $\rho:A\rightarrow A$ such that $\rho_{2}(ab)=a\rho_{2}(b)$ for all $a,b\in A$. A multiplier of AA consists of a pair $(\rho_{1},\rho_{2})$ comprising a left and a right multiplier, where $\rho_{2}(a)b=a\rho_{1}(b)$ holds for all $a,b\in A$. For any algebra AA, there exist a left multiplier algebra, a right multiplier algebra, and a multiplier algebra of $A$, denoted as $L(A)$, $R(A)$, and $M(A)$, respectively.
Specifically, $M(A)$ is characterized as the largest algebra with identity containing AA as an essential two-sided ideal. Moreover, in the case where $x\in M(A)$ and $xa=0$ for all $a\in A$ or $ax=0$ for all $a\in A$, it follows that $x=0$. Additionally, we examine the tensor algebra $A\otimes A$, which remains nondegenerate, and we have its multiplier algebra $M(A\otimes A)$. There are natural embeddings
@@
Generally, when AA lacks an identity element, these two embeddings are strict. In the scenario where AA possesses an identity element $1_{A}$, the product is evidently nondegenerate, and we observe that $M(A)=A$ and $M(A\otimes A)=A\otimes A$.
Let $A$ and $B$ be nondegenerate algebras. If a homomorphism $f:A\longrightarrow M(B)$ is nondegenerate (meaning $f(A)B=B$ and $Bf(A)=B$), then it can be uniquely extended to a homomorphism $M(A)\longrightarrow M(B)$, also denoted by $f$.
A multiplier Hopf coquasigroup, as defined in [14], is a nondegenerate associative algebra AA equipped with algebra homomorphisms $\Delta:A\longrightarrow M(A\otimes A)$ (comultiplication), $\varepsilon:A\longrightarrow k$ (counit), and a linear map $S:A\longrightarrow A$ (antipode). The key conditions for this structure are as follows:
@@
    For any $a,b\in A$, both $T_{1}(a\otimes b)=\Delta(a)(1\otimes b)$ and $T_{2}(a\otimes b)=(a\otimes 1)\Delta(b)$ belong to $A\otimes A$.
@@
    The counit satisfies $(\varepsilon\otimes\iota)T_{1}(a\otimes b)=ab=(\iota\otimes\varepsilon)T_{2}(a\otimes b)$, where $\iota$ denotes the identity map, and we will use this notation henceforth.
@@
    The antipode $S$ is defined to be antimultiplicative and anticomultiplicative, such that for any $a,b\in A$ the following hold:
@@
In the context of a multiplier Hopf coquasigroup $(A,\Delta)$, it is termed regular if the antipode $S$ is bijective. The maps $T_{1}$ and $T_{2}$ are often called the Galois maps.
Multiplier Hopf coquasigroups, as introduced in [14], extend the concept of Hopf coquasigroups to a non-unital scenario. In contrast to the Hopf coquasigroups discussed in [5, 6], the algebraic structure of a multiplier Hopf coquasigroup AA may not necessarily possess an identity element in general. Moreover, the comultiplication map $\Delta$ does not merely map elements from $A$ to $A\otimes A$, but rather it maps them into the multiplier algebra $M(A\otimes A)$. This extension weakens the coassociativity property typically found in multiplier Hopf algebras while also harmonizing the concepts of multiplier Hopf algebras and Hopf coquasigroups.
To denote the comultiplication, we employ the adapted Sweedler notation as introduced in [12]: $\Delta(a)=a_{(1)}\otimes a_{(2)}$ for $a\in A$.
2.2 Ore extension
@@
We start by recalling the definition of the Ore extension of an algebra, as introduced in [4]. Given an algebra AA with an identity element $1_{A}$ and an algebra endomorphism $\tau$ of $A$, a linear endomorphism $\delta$ of $A$ is referred to as a $\tau$-derivation if for all $a,b\in A$,
$\delta(ab)=\delta(a)b+\tau(a)\delta(b).$
Moreover, it is a consequence of the above condition that $\delta(1_{A})=0$.
The Ore extension $A[y;\tau,\delta]$ of an algebra AA is constructed as an algebra that extends AA with the defining relation
@@
Ore extension of Hopf algebras
Let $A$ and $A[y;\tau,\delta]$ denote Hopf algebras over $k$. The Hopf algebra $A[y;\tau,\delta]$ is termed a Hopf-Ore extension if there exists an element $r\in A$ such that $\Delta(y)=y\otimes 1+r\otimes y$, and $A$ is a Hopf subalgebra of $A[y;\tau,\delta]$.
In accordance with Theorem 1.3 in [9], the conditions necessary and sufficient for the Hopf algebra $A[y;\tau,\delta]$ to qualify as a Hopf-Ore extension are as follows:
$(1)$ There exists a character $\chi:A\rightarrow k$ satisfying
@@
$(2)$ The following condition holds
@@
$(3)$ The $\tau$-derivation $\delta$ fulfills the relation
@@
Ore extension of Multiplier Hopf algebras
By generalizing the Ore extension of Hopf algebras to the non-unital case, Zhao and Lu [16] introduced the Ore extension of multiplier Hopf algebras.
For more details, let $(A,\Delta)$ be a regular multiplier Hopf algebra, and consider $A[y;\tau,\delta]$, an Ore extension of the algebra AA. When $\tau$ is a nondegenerate algebra endomorphism, the $\tau$-derivation $\delta$ can be uniquely extended to the multiplier algebra $M(A)$, denoted as $\overline{\delta}:M(A)\rightarrow M(A)$. In this case,
$A[y;\tau,\delta]\subseteq M(A)[y;\tau,\delta]\subseteq M(A[y;\tau,\delta]).$
Where $\tau$ is surjective, according to Proposition 2.3 of Section 2 in [16], the product of $A[y;\tau,\delta]$ is nondegenerate.
Zhao and Lu first constructed $M(A)[y;\tau,\delta]$, an Ore extension of $M(A)$. The comultiplication, counit, and antipode of AA can be extended to $M(A)[y;\tau,\delta]$: $\overline{\Delta}:M(A)[y;\tau,\delta]\rightarrow M(A[y;\tau,\delta]\otimes A[y;\tau,\delta])$, $\overline{\varepsilon}:M(A)[y;\tau,\delta]\rightarrow k$, $\overline{S}:M(A)[y;\tau,\delta]\rightarrow M(A[y;\tau,\delta])$. By restricting these extensions to the subspace $A[y;\tau,\delta]$, the authors obtained the comultiplication, counit, and antipode on the Ore extension $A[y;\tau,\delta]$ of AA:
$\overline{\Delta}\big{|}_{A[y;\tau,\delta]}:A[y;\tau,\delta]\rightarrow M(A[y;\tau,\delta]\otimes A[y;\tau,\delta]),$
$\overline{\varepsilon}\big{|}_{A[y;\tau,\delta]}:A[y;\tau,\delta]\rightarrow k,$
$\overline{S}\big{|}_{A[y;\tau,\delta]}:A[y;\tau,\delta]\rightarrow M(A[y;\tau,\delta]).$
These maps were respectively denoted by $\Delta^{\prime}$, $\varepsilon^{\prime}$, and $S^{\prime}$.
The pair $(A[y;\tau,\delta],\Delta^{\prime})$ is called a Hopf-Ore extension of the multiplier Hopf algebra AA, if $(A[y;\tau,\delta],\Delta^{\prime},\varepsilon^{\prime},S^{\prime})$ is a regular multiplier Hopf algebra such that $\overline{\Delta}(y)=y\otimes 1_{M(A)}+r\otimes y$ for some $r\in M(A)$, where $1_{M(A)}$ is the identity element of $M(A)$.
For a regular multiplier Hopf algebra AA with an Ore extension $A[y;\tau,\delta]$, where $\tau$ is surjective, $A[y;\tau,\delta]$ forms a Hopf-Ore extension if and only if there exists a group-like element $r\in M(A)$ satisfying the following equivalent conditions:
$\tau(a)=(\chi\otimes\iota)\Delta(a)=r\big{(}(\iota\otimes\chi)\Delta(a)\big{)}r^{-1}.$
$(2)$ For the $\tau$-derivation $\delta$, for all $a\in A$, the relation
$\Delta\big{(}\delta(a)\big{)}=(\delta\otimes\iota)\Delta(a)+(r\otimes 1_{M(A)})\big{(}(\iota\otimes\delta)\Delta(a)\big{)}$
is satisfied.
As described in Section 2.1 of this paper, we will refer to these unique extensions as $\overline{\Delta}$, $\overline{S}$, $\overline{\varepsilon}$ simply as $\Delta$, $S$, $\varepsilon$ in the following without causing any confusion.
@@
Let $A$ and $A[y;\tau,\delta]$ denote Hopf coquasigroups. As defined in [3], the Hopf coquasigroup $A[y;\tau,\delta]$ is classified as a Hopf coquasigroups-Ore extension when $\Delta(y)=y\otimes 1+r\otimes y$ for a certain $r\in A$, and $A$ is identified as a sub-Hopf coquasigroup of $A[y;\tau,\delta]$.
According to Theorem 3.3 in [3], a Hopf coquasigroup AA transforms into a Hopf coquasigroup-Ore extension denoted by $H=A[y,\tau,\delta]$ if and only if the following conditions are met:
$(1)$ A character $\chi:A\rightarrow k$ exists such that for each $a\in A$, the condition $\tau(a)=\chi(a_{(1)})a_{(2)}$ is satisfied.
(2)(2)The subsequent relations are valid:
@@
$(3)$ For the $\tau$-derivation $\delta$, the relation
@@ 3 Ore extensions of multiplier Hopf coquasigroups
Let $(A,\Delta)$ denote a regular multiplier Hopf coquasigroup as defined in [14]. The objective of this section is to establish a criterion for an Ore extension of a regular multiplier Hopf coquasigroup to also qualify as a regular multiplier Hopf coquasigroup.
@@
First, we shall introduce the definition of the Ore extension of regular multiplier Hopf coquasigroups, which extends the Ore extensions of Hopf coquasigroups and multiplier Hopf algebras as presented in [3, 16].
@@
Definition 3.1. Let $A$ and $A[y;\tau,\delta]$ be regular multiplier Hopf coquasigroups. The multiplier Hopf coquasigroup $(A[y;\tau,\delta],\Delta)$ is called an Ore extension of multiplier Hopf coquasigroup (or an MHC-Ore extension in short) if the following equation holds:
$\Delta(y)=y\otimes 1_{M(A)}+r\otimes y$ (3.1)
for a group-like element $r\in M(A)$.
Remark 3.2. $(1)$ In what follows, we will simply write $1_{M(A)}$ as $1$ without causing any confusion. As stated in [16], the element $y\otimes 1+r\otimes y$ in equation $(\ref{3.1})$ belongs to $M(A[y;\tau,\delta]\otimes A[y;\tau,\delta])$. As usual, we denote $Ad_{r}(a)=raS(r)=rar^{-1}$.
$(2)$ When the comultiplication $\Delta$ of the multiplier Hopf coquasigroup $A$ satisfies coassociativity, $A$ degenerates to a multiplier Hopf algebra and the Ore extension aligns with the one outlined in [16]. As introduced in Section 2 of [16], all extensions depend on the algebraic structure of $A$, then extensions on multiplier Hopf coquasigroups are similar.
@@
$(3)$ If the underlying algebra of the multiplier Hopf coquasigroup AA possesses an identity element, $A$ can be classified as a Hopf coquasigroup, and the Ore extension corresponds to the one delineated in [3].
@@
Now, we present the properties associated with the Ore extension of a regular multiplier Hopf coquasigroup.
Proposition 3.3. When $A[y;\tau,\delta]$ is an MHC-Ore extension of $A$, then
$\varepsilon(y)=0,\quad S(y)=-r^{-1}y.$
Proof. By applying the counit map $\varepsilon\otimes\iota$ to both sides of equation $(\ref{3.1})$, we obtain
$y$ $=(\varepsilon\otimes\iota)\Delta(y)$
$=(\varepsilon\otimes\iota)(y\otimes 1+r\otimes y)$
$=\varepsilon(y)1+\varepsilon(r)y$
$=\varepsilon(y)1+y.$
Hence, it follows that $\varepsilon(y)=0$.
Next, applying the same map $\varepsilon\otimes\iota$ again to equation $(\ref{3.1})$, we have
$\varepsilon(y)$ $=m(\iota\otimes S)\Delta(y)$
$=m(\iota\otimes S)(y\otimes 1+r\otimes y)$
$=y+rS(y).$